# A Collision-Free Hot-Tier Extension for Engram-Style Conditional Memory: A Controlled Study of Training Dynamics

### v1

@@ 1.1 Background and Motivation
The parameter scale of large language models (LLMs) continues to grow, yet research has shown that a considerable portion of model capacity is devoted to memorizing “local, stereotyped, entity/phrase-style” static patterns [1]. The Engram module proposed by DeepSeek offloads these static patterns from the Transformer’s attention mechanism to an external memory module via O(1) n-gram hash lookup, thereby freeing backbone capacity for more complex reasoning tasks. The appeal of this approach lies in its deterministic indexing: retrieval indices depend solely on the input token sequence, enabling index pre-computation and embedding prefetching on the host side during inference, offloading of embedding tables to host memory or NVMe storage, and prefetching of multiple candidate paths during speculative decoding.
However, the original Engram’s multi-head hashing mechanism has an apparent “weakness”: hash collisions. When two semantically different n-grams map to the same slot, they are forced to share the same embedding row, effectively injecting incorrect prior information into the model. Since n-gram access follows a Zipf distribution [5]—a small number of high-frequency keys account for the majority of queries—the “total damage” from collisions is greatest for high-frequency keys. A natural intuition is: eliminating collisions for high-frequency keys should lead to lower loss.
@@ 1.2 Main Contributions
We report a counterintuitive but reproducible result: under strict iso-parameter (parameter-equivalent) control, replacing the high-frequency n-gram index from “high-collision hashing” to “collision-free MPHF (hot tier)” does not improve validation loss. This result suggests that in Engram-style hashed memory, high-frequency collisions are not necessarily the dominant bottleneck; optimizing index precision alone is insufficient to guarantee training benefits.
To explain this phenomenon, we propose a route-stratified evaluation: on the same model checkpoint, we decompose each position’s loss by hot/cold path, combined with gate weight $\alpha$ bucketing and ablation experiments to analyze training dynamics, thereby advancing the question from “is there benefit” to a mechanistic diagnosis of “why no benefit.”
The main contributions of this paper include: (1) the first iso-parameter comparison between collision-prone and collision-free variants of Engram, providing rigorous controlled experimental evidence; (2) a diagnostic methodology combining hot/cold loss decomposition with alpha bucket analysis, offering methodological reference for debugging similar systems; (3) mechanism-level discoveries including the hot/cold flip phenomenon, implicit regularization effects of collisions, and gating mismatch issues; (4) a warning to the community: improving retrieval precision does not equal improving training benefits—deeper mechanistic understanding is needed.
@@ 2.1 External Memory and Retrieval Augmentation
Neural network external memory mechanisms have a long research history. Early Neural Turing Machines [6] and Memory Networks [7] explored differentiable read-write mechanisms, but their soft attention mechanisms have high computational costs at scale. Product Key Memory [8] achieves O($\sqrt{N}$) approximate retrieval through product quantization, demonstrating the value of external memory in language modeling. In recent years, Retrieval-Augmented Generation (RAG) series works [9, 10] enhance generation capability by retrieving external documents, but their retrieval granularity is at the document or passage level, fundamentally different from token-level n-gram memory in design goals.
@@ 2.2 Hashing and Perfect Hashing
Hash collisions are an inherent problem of traditional hash tables. Cuckoo Hashing [11] achieves O(1) worst-case query through multi-table probing, but has lower space efficiency (load factor $\sim$0.5). Minimal Perfect Hash Functions (MPHF) provide collision-free mapping [0,N1][0,N-1] for static key sets. In recent years, tools like BBHash [2] and PTHash [12] have made MPHF construction efficient and practical (approximately 2–4 bits/key, depending on algorithm and parameters). We chose MPHF as the indexing scheme for the hot tier precisely because of its compactness and determinism.
@@ 2.3 Regularization and Noise Injection
Regularization techniques in machine learning—Dropout [13], Label Smoothing [14], data augmentation, etc.—prevent overfitting by introducing controlled noise. Our experimental results suggest that hash collisions may play a similar role: they force semantically similar n-grams to share embeddings, equivalent to a form of implicit clustering or averaging. This observation echoes the findings of Hash Embeddings [15], which noted that despite information loss from hash collisions, the method “often works surprisingly well” in practice.
@@ 3.2 Engram-Nine’s Two-Tier Retrieval Architecture
Hot path: Uses MPHF to obtain a unique index idx, retrieving a complete $K\times d$ dimensional vector from a single embedding table $E_{\text{hot}}\in\mathbb{R}^{N_{\text{hot}}\times Kd}$. Since MPHF is collision-free for the hot key set, each hot key has its own dedicated row.
Cold path: Follows the original Engram’s multi-head hashing. For the same n-gram, use $K$ different hash functions $\phi_{1},\ldots,\phi_{K}$ to compute indices separately, retrieving one $d$-dimensional vector from each of $K$ independent tables $E^{(k)}_{\text{cold}}\in\mathbb{R}^{M\times d}$, concatenating to obtain $K\times d$ dimensional output. Due to hash collisions, different n-grams may share the same row.
@@ 4.1 Training Scale and Data
Backbone GPT-2 architecture (125M params, 12 layers, 768 dim)
@@
Training uses the nanoGPT [3] framework with AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay = 0.1). Learning rate is set to $6\times 10^{-4}$ with cosine annealing schedule (warmup 100 steps, decay to $6\times 10^{-5}$). Batch size is 4, gradient accumulation 4 steps, effective batch size 16. Total training is 5000 steps (approximately 82M tokens). Each configuration is run with 2 different random seeds to assess variance. Hardware is a single NVIDIA A100-SXM4-40GB GPU.
@@ 4.3 Evaluation Metrics
Route Labeling (is_hot mask)
At each evaluation, for each token position, we label whether its n-gram belongs to the top-$N_{\text{hot}}$ high-frequency set. The purpose of this label is to answer: “If this position were in Engram-Nine, would it take the hot path or cold path?”
It should be emphasized that the Hash baseline does not have native hot/cold routing during training—all n-grams take the same multi-head hash path. Therefore, the Hash baseline’s is_hot label is a post-hoc annotation, purely for analysis purposes: it tells us, for the same batch of validation data, which positions would be routed to the hot tier in the Nine version. By using the same top-$N_{\text{hot}}$ frequency threshold for labeling both configurations, we can fairly compare: “Under the same hot/cold division, how do the collision-prone Hash and collision-free Nine each perform?”
@@
Based on the is_hot mask, we compute separately:
  • val_loss_hot: Average loss for hot positions only (is_hot=True)
  • val_loss_cold: Average loss for cold positions only (is_hot=False)
@@
Similar to stratified loss, we compute gating weights by is_hot mask:
  • $\alpha_{\text{hot}}$: Average gating weight for hot positions (is_hot=True)
  • $\alpha_{\text{cold}}$: Average gating weight for cold positions (is_hot=False)
@@ 5 Experimental Results
5.1 Main Conclusion: Eliminating Collisions Yields No Significant Benefit
@@ 5.2 Core Finding: Hot/Cold Flip Phenomenon
Figure 2 shows the evolution of hot_cold_delta ($\text{val\_loss\_hot}-\text{val\_loss\_cold}$) over training steps. In early training (iter 1000–2000), hot positions’ loss is notably lower than cold positions (delta $\approx-0.15\sim-0.22$), consistent with the intuition that “hot keys are easier to predict”—high-frequency n-grams have more concentrated successor distributions, making them easier for the model to learn. However, by late training (iter 4000+), cold positions overtake, with the gap widening to $0.07\sim+0.17$.
[…]
@@ 5.3 $N_{\text{hot}}$ Ablation: Refuting the Sparsity Hypothesis
This counterintuitive result has two possible explanations: (1) Under iso-parameter design, smaller $N_{\text{hot}}$ means larger cold_table_size, lower collision rate in the cold tier—and we’ve already shown collisions have a regularization effect that delays the flip, so the cold tier with fewer collisions “learns well” earlier and overtakes. However, this explanation has a contradiction: pure Hash configurations are equivalent to 100% cold tier, yet their flip time is iter 3000, later than all Nine configurations. This indicates what truly delays the flip is not the cold table’s collision rate level, but the hot tier’s collision-free property itself—Nine’s hot portion lacks regularization and is more prone to overfitting. (2) Smaller $N_{\text{hot}}$ means lower hot tier coverage (41% vs 59%), hot’s influence on overall loss is inherently smaller, and the flip signal is more easily amplified. Regardless of explanation, both indicate the root cause is not “insufficient training samples for hot embeddings.”
@@ 5.4 Gating Preference Analysis
5.4.1 Evolution of $\alpha_{\text{hot}}$ vs $\alpha_{\text{cold}}$ During Training
Table 6 shows the changes in $\alpha_{\text{hot}}$ and $\alpha_{\text{cold}}$ over training iterations for the Nine-100K configuration, along with the concurrent hot/cold loss difference.
Table 6: Evolution of $\alpha_{\text{hot}}$ and $\alpha_{\text{cold}}$ During Training (Nine-100K, seed0)
Iter $\alpha_{\text{hot}}$ $\alpha_{\text{cold}}$ $\Delta\alpha$ $_{\text{hot}} - Loss$$_{\text{cold}}$ Preference vs Performance
1000 0.78 0.69 +0.09+0.09 0.19-0.19 Aligned (hot better, prefers hot)
2000 0.72 0.63 +0.09+0.09 0.09-0.09 Aligned
3000 0.69 0.61 +0.08+0.08 $\approx 0$ Critical point
4000 0.66 0.60 +0.06+0.06 +0.04+0.04 Diverged (cold better, still prefers hot)
4750 0.66 0.61 +0.05+0.05 +0.09+0.09 Divergence intensified
This data reveals several key patterns:
Pattern 1: Gate Consistently Prefers Hot
From iter 1000 to iter 4750, $\alpha_{\text{hot}}>\alpha_{\text{cold}}$ always holds, indicating the gate learned to distinguish hot/cold and formed a “hot keys are more trustworthy” preference. This pattern stably reproduces in both seed0 and seed2.
Pattern 2: Preference and Performance Diverge After Iter 3000
[…]
Pattern 3: Overall $\alpha$ Decline
Both hot and cold $\alpha$ decline from early 0.7–0.8 to late 0.6–0.7. This indicates the model overall becomes more “cautious,” with reduced trust in memory injection. But the relative relationship $\alpha_{\text{hot}}>\alpha_{\text{cold}}$ is always maintained.
5.4.2 Static Analysis by $\alpha$ Bucketing
@@
The data clearly confirms the dynamic analysis conclusion: the model assigns higher $\alpha$ to positions with higher loss. Low $\alpha$ (0–0.4) positions have loss around 3.9, while high $\alpha$ (0.8–1.0) positions have loss as high as 5.1–5.3—completely opposite to the gating design intent. More critically, approximately 70% of the high $\alpha$ bucket are hot positions (Hash 76%, Nine 69%), corroborating the finding that $\alpha_{\text{hot}}>\alpha_{\text{cold}}$”: gate prefers hot, but hot’s loss is actually higher in late training.
@@ 5.4.3 Synthesis: Preference Fixation Hypothesis
Synthesizing dynamic and static analyses, we propose an explanation: gating preference forms and fixates in early training. The specific mechanism is as follows:
@@ 5.5 Layer-Level Anomaly: Layer 6’s $\alpha$ Reversal
  • Layer 4 (middle): Most configurations still prefer hot tier, but Nine-100K begins showing slight cold preference.
@@ 6.1 Precision $\neq$ Better: A More General Principle
Our experiments clearly demonstrate: eliminating collisions does not equal better training benefits. Under iso-parameter conditions, Engram-Nine’s collision-free hot tier did not bring the expected loss improvement, but instead led to earlier and more severe hot/cold flipping. This result challenges the intuitive assumption that “collisions are the main bottleneck.”

### v2

@@ 1.1 Background and Motivation
The parameter scale of large language models (LLMs) continues to grow, yet research has shown that a considerable portion of model capacity is devoted to memorizing “local, static, and highly stereotyped” patterns [1]. The Engram module proposed by DeepSeek [1] offloads these static patterns from the Transformer’s attention mechanism to an external memory module via O(1) n-gram hash lookup, thereby freeing backbone capacity for more complex reasoning tasks. The appeal of this approach lies in its deterministic indexing: retrieval indices depend solely on the input token sequence, enabling index pre-computation and embedding prefetching on the host side during inference, offloading of embedding tables to host memory or NVMe storage, and prefetching of multiple candidate paths during speculative decoding.
However, the original Engram’s multi-head hashing mechanism has an apparent “weakness”: hash collisions. When two semantically different n-grams map to the same slot, they are forced to share the same embedding row, effectively injecting incorrect prior information into the model. Since n-gram access follows a Zipf distribution [2]—a small number of high-frequency keys account for the majority of queries—collisions on high-frequency keys are encountered far more often, amplifying their cumulative impact on the model. A natural intuition is: eliminating collisions for high-frequency keys should lead to lower loss.
@@ 1.2 Main Contributions
We report a counterintuitive but reproducible result: under strict iso-parameter (i.e., both methods use exactly the same number of trainable parameters) control, replacing the high-frequency n-gram index from “high-collision hashing” to “collision-free MPHF (hot tier)” does not improve validation loss. This result suggests that in Engram-style hashed memory, high-frequency collisions are not necessarily the dominant bottleneck; optimizing index precision alone is insufficient to guarantee training benefits.
To explain this phenomenon, we propose a route-stratified evaluation: on the same model checkpoint, we decompose each position’s loss by hot/cold tier, combined with gate weight $\alpha$ bucketing and ablation experiments to analyze training dynamics, thereby advancing the question from “is there benefit” to a mechanistic diagnosis of “why no benefit.”
The main contributions of this paper include: (1) the first iso-parameter comparison between collision-prone and collision-free variants of Engram, providing rigorous controlled experimental evidence; (2) a diagnostic methodology combining hot/cold loss decomposition with alpha bucket analysis, offering methodological reference for debugging similar systems; (3) mechanism-level discoveries including the hot/cold flip phenomenon, implicit regularization effects of collisions, and gating mismatch issues; (4) empirical evidence cautioning that improving retrieval precision does not necessarily improve training benefits—deeper mechanistic understanding is needed.
@@ 2.1 External Memory and Retrieval Augmentation
Neural network external memory mechanisms have a long research history. Early Neural Turing Machines [3] and Memory Networks [4] explored differentiable read-write mechanisms, but their soft attention mechanisms have high computational costs at scale. Product Key Memory [5] achieves O($\sqrt{N}$) approximate retrieval through product quantization, demonstrating the value of external memory in language modeling. In recent years, Retrieval-Augmented Generation (RAG) series works [6, 7] enhance generation capability by retrieving external documents. A key distinction is that RAG operates at inference time—dynamically retrieving from external corpora during generation—while Engram integrates memory at training time, learning to utilize n-gram statistics as the model trains. Additionally, RAG’s retrieval granularity is at the document or passage level, whereas Engram operates at token-level n-gram precision.
@@ 2.2 Hashing and Perfect Hashing
Hash collisions are an inherent problem of traditional hash tables. Cuckoo Hashing [8] achieves O(1) worst-case query through multi-table probing, but has lower space efficiency (load factor $\sim$0.5). Minimal Perfect Hash Functions (MPHF) [9] provide collision-free mapping [0,N1][0,N-1] for static key sets—that is, $N$ keys are mapped to exactly $N$ consecutive slots with no collisions and no wasted space, achieving the theoretical minimum storage for a bijective hash. In recent years, tools like BBHash [10] and PTHash [11] have made MPHF construction efficient and practical (approximately 2–4 bits/key, depending on algorithm and parameters). We chose MPHF as the indexing scheme for the hot tier precisely because of its compactness and determinism.
@@ 2.3 Regularization and Noise Injection
Regularization techniques in machine learning—Dropout [12], Label Smoothing [13], data augmentation, etc.—prevent overfitting by introducing controlled noise. Our experimental results suggest that hash collisions may play a similar role: they force semantically similar n-grams to share embeddings, equivalent to a form of implicit clustering or averaging. This observation echoes the findings of Hash Embeddings [14], which achieves comparable performance to traditional embeddings despite hash collisions causing different tokens to share embedding vectors.
@@ 3.2 Engram-Nine’s Two-Tier Retrieval Architecture
Hot tier: Uses MPHF to obtain a unique index idx, retrieving a complete $K\times d$ dimensional vector from a single embedding table $E_{\text{hot}}\in\mathbb{R}^{N_{\text{hot}}\times Kd}$. Since MPHF is collision-free for the hot key set, each hot key has its own dedicated row.
Cold tier: Follows the original Engram’s multi-head hashing. For the same n-gram, use $K$ different hash functions $\phi_{1},\ldots,\phi_{K}$ to compute indices separately, retrieving one $d$-dimensional vector from each of $K$ independent tables $E^{(k)}_{\text{cold}}\in\mathbb{R}^{M\times d}$, concatenating to obtain $K\times d$ dimensional output. Due to hash collisions, different n-grams may share the same row.
@@ 4.1 Training Scale and Data
Backbone GPT-2 architecture ($\sim$185M params, 12 layers, 768 dim)
@@
Larger than standard GPT-2 (124M) due to expanded vocabulary (128,815 vs 50,257).
Parameter Sharing Across Layers
The three Engram layers (Layer 2, 4, 6) share a single set of n-gram embedding tables to avoid parameter explosion. However, each layer maintains its own independent gating parameters ($W_{Q}$, $W_{K}$, $W_{V}$, RMSNorm), depthwise convolution, and output projection. This design keeps the model compact while allowing layer-specific modulation—as evidenced by the per-layer gating differences analyzed in Section 5.4.
@@
Training uses the nanoGPT [16] framework with AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay = 0.1). Learning rate is set to $6\times 10^{-4}$ with cosine annealing schedule (warmup 100 steps, decay to $6\times 10^{-5}$). Batch size is 4, gradient accumulation 4 steps, effective batch size 16. Total training is 5000 steps (approximately 82M tokens). Each configuration is run with 3 different random seeds to assess variance. Hardware is a single NVIDIA A100-SXM4-40GB GPU.
@@ 4.2 Iso-Parameter Comparison Design
Total = GPT-2 backbone ($\sim$185M) + shared Engram embeddings (128M) + per-layer gating/conv/projection ($\sim$1.6M for 3 layers).
Position embeddings excluded following common practice.
@@ 4.3 Evaluation Metrics
Route Labeling (Hot-Tier Mask)
At each evaluation, for each token position, we label whether its n-gram belongs to the top-$N_{\text{hot}}$ high-frequency set. The purpose of this label is to answer: “If this position were in Engram-Nine, would it be routed to the hot tier or cold tier?”
It should be emphasized that the Hash baseline does not have native hot/cold routing during training—all n-grams take the same multi-head hash path. Therefore, the Hash baseline’s hot-tier label is a post-hoc annotation, purely for analysis purposes: it tells us, for the same batch of validation data, which positions would be routed to the hot tier in the Nine version. By using the same top-$N_{\text{hot}}$ frequency threshold for labeling both configurations, we can fairly compare: “Under the same hot/cold division, how do the collision-prone Hash and collision-free Nine each perform?”
@@
Based on the hot-tier mask, we compute separately:
  • val_loss_hot: Average loss for hot-tier positions
  • val_loss_cold: Average loss for cold-tier positions
@@
Similar to stratified loss, we compute gating weights by the hot-tier mask:
  • $\alpha_{\text{hot}}$: Average gating weight for hot-tier positions
  • $\alpha_{\text{cold}}$: Average gating weight for cold-tier positions
@@ 5 Experimental Results
5.1 Main Finding: Eliminating Collisions Yields No Significant Benefit
@@ 5.2 Core Finding: Hot/Cold Flip Phenomenon
Figure 2 shows the evolution of hot_cold_delta ($\text{val\_loss\_cold}-\text{val\_loss\_hot}$) over training steps. In early training (iter 1000–2000), hot positions’ loss is notably lower than cold positions (delta $\approx+0.15\sim+0.22$), consistent with the intuition that “hot keys are easier to predict”—high-frequency n-grams have more concentrated successor distributions, making them easier for the model to learn. However, by late training (iter 4000+), cold positions overtake, with the gap widening to $0.07\sim-0.17$.
Figure 2: Training dynamics of the Hot/Cold flip phenomenon. All configurations transition from early hot advantage (delta >> 0) to late cold advantage (delta << 0). Nine configurations flip earlier than Hash configurations. (Hash-500K only preserved the final checkpoint, so it is not included in flip timing analysis.)
@@ 5.3 $N_{\text{hot}}$ Ablation: Refuting the Sparsity Hypothesis
The results directly contradict the sparsity hypothesis: the configuration with the densest samples per embedding (Nine-20K) flips earliest rather than latest. This rules out “insufficient training samples for hot embeddings” as the root cause.
What alternative mechanisms might explain the observed pattern? We consider two possibilities:
Hypothesis A: Cold-tier collision rate.
Under iso-parameter design, smaller $N_{\text{hot}}$ yields a larger cold table with lower collision rates. Since collisions provide implicit regularization that delays the flip (Section 6.2), a less-collided cold tier might “learn well” earlier and overtake sooner. However, this explanation faces a difficulty: pure Hash configurations—effectively 100% cold tier—flip at iter 3000, later than all Nine configurations. This suggests the flip timing may be more strongly influenced by the presence of a collision-free hot tier than by the cold table’s collision rate alone, though disentangling these factors would require additional controlled experiments (e.g., artificially injecting collisions into the hot tier).
Hypothesis B: Hot-tier coverage.
Smaller $N_{\text{hot}}$ implies lower hot-tier coverage (41% vs. 59%), meaning the hot tier contributes less to overall loss. Consequently, even a modest degradation in hot-tier performance produces a more pronounced flip signal in the aggregate metrics.
Both hypotheses attribute the flip timing to architectural factors (collision structure, coverage ratio) rather than to per-embedding sample density, consistent with our rejection of the sparsity hypothesis.
@@ 5.4 Gating Preference Analysis
5.4.1 Temporal Analysis: Evolution of $\alpha$ During Training
Figure 3 shows the evolution of raw gate weights ($\alpha_{\text{hot}}$, $\alpha_{\text{cold}}$), gate preference ($\Delta\alpha=\alpha_{\text{hot}}-\alpha_{\text{cold}}$), and actual performance difference ($_{\text{cold}} - Loss$$_{\text{hot}}$) over training iterations for Nine-100/400K.
Figure 3: Gate weight evolution during training (Nine-100/400K, seed0). Top: Raw $\alpha$ values show overall decline from $\sim$1.0 to $\sim$0.65. Middle: Gate preference $\Delta\alpha$ stays positive throughout, indicating persistent hot preference. Bottom: After iter 2500, cold achieves lower loss (negative region), but gate preference does not reverse—the preference “crystallized” in early training.
This visualization reveals several key patterns:
Pattern 1: Overall $\alpha$ Decline
Both hot and cold $\alpha$ decline from early $\sim$1.0 to late 0.6–0.7. This indicates the model overall becomes more “cautious,” with reduced trust in memory injection.
Pattern 2: Gate Consistently Prefers Hot
Despite the overall decline, $\alpha_{\text{hot}}>\alpha_{\text{cold}}$ always holds from iter 1000 to iter 4750, indicating the gate learned to distinguish hot/cold and formed a “hot keys are more trustworthy” preference. This pattern stably reproduces across all seeds.
Pattern 3: $\Delta\alpha$ Gradually Shrinking
While the gate maintains hot preference throughout training, the preference gap $\Delta\alpha=\alpha_{\text{hot}}-\alpha_{\text{cold}}$ is gradually shrinking. This trend is consistent across all three seeds and shows no sign of plateauing by iter 4750. This suggests that given sufficient training time, the gate might eventually correct its hot bias—but the current training duration is insufficient to observe this potential correction.
Pattern 4: Preference and Performance Diverge
In early training (iter 1000–2000), the gate’s preference aligns with actual performance: hot loss is indeed lower, gate gives hot higher $\alpha$. However, from iter 2500 onward, cold loss becomes lower than hot loss, but the gate’s preference does not reverse. In other words, the gate’s preference “crystallized” once formed in early training, creating a mismatch between gating behavior and actual prediction performance.
5.4.2 Snapshot Analysis: $\alpha$ Bucketing
@@
The data clearly confirms the temporal analysis conclusion: the model assigns higher $\alpha$ to positions with higher loss. Low $\alpha$ (0–0.4) positions have loss around 3.9, while high $\alpha$ (0.8–1.0) positions have loss as high as 5.1–5.3—completely opposite to the gating design intent. More critically, approximately 70% of the high $\alpha$ bucket are hot positions (Hash 76%, Nine 69%), corroborating the finding that $\alpha_{\text{hot}}>\alpha_{\text{cold}}$”: gate prefers hot, but hot’s loss is actually higher in late training.
@@ 5.4.3 Synthesis: Preference Fixation Hypothesis
Synthesizing temporal and snapshot analyses, we propose an explanation: gating preference forms and fixates in early training. The specific mechanism is as follows:
@@ 5.5 Layer-Level Anomaly: Layer 6’s $\alpha$ Reversal
  • Layer 4 (middle): Most configurations still prefer hot tier, but Nine-100/400K begins showing slight cold preference.
@@ 6.1 Precision $\neq$ Better: A More General Principle
Our experiments suggest: at this scale, eliminating collisions does not equal better training benefits. Under iso-parameter conditions, Engram-Nine’s collision-free hot tier did not bring the expected loss improvement, but instead led to earlier and more severe hot/cold flipping. This result challenges the intuitive assumption that “collisions are the main bottleneck.”